An entrance test is made up of three parts. Each part has 20 questions. What is the maximum number of questions one can fail and still get an overall average of 70% to pass the test? pls, show full working so i can understand it. Thanks.
Answer by stanbon(75887)

An entrance test is made up of three parts. Each part has 20 questions. What is the maximum number of questions one can fail and still get an overall average of 70% to pass the test? pls, show full working so i can understand it.
--------
There are a total of 3*20 = 60 questions.
---
To get a grade of 70% you need to get 0.7*60 = 42 questions correct.
----
The maximum number you can fail is 60-42 = 18.
-----------
Cheers,
Stan H.